OpenStudy (anonymous):

Determine if (1/2,-4) is a solution to 4x-3y=-10. Please show all work.

OpenStudy (anonymous):

4(1/2)-3(-4)=-10 2+12=-10 False

OpenStudy (gw2011):

4x-3y=-10 Substitute the values into the equation (4)(1/2)-(3)(-4)=2+12=14 Therefore, (1/2, -4) is not a solution to gthe equation.